OpenStudy (anonymous):

will give metal to best answer - which measure id the greatest? A. 7 kg B. 68.1 cg C. 700 mg D. 6,900 g

OpenStudy (anonymous):

Convert to similar units and then it'll be easy

OpenStudy (anonymous):

umm

OpenStudy (studygurl14):

kg = 1000g cg = g/100 mg = g/1000 There's the info you need

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

i still don't know what i am doing....

OpenStudy (anonymous):

1 kg = 1000 grams 7 kg = ? grams

OpenStudy (anonymous):

1 centigram = 0.01 grams 68.1 cg = ?? gms

OpenStudy (anonymous):

Another way 1/1000 = 7 /x solve for x

OpenStudy (anonymous):

1/0.01 = 68.1/x

OpenStudy (anonymous):

B

OpenStudy (anonymous):

1 gram = 1000 milligrams 1/1000 = x/700

OpenStudy (anonymous):

howd you get that answer?

OpenStudy (anonymous):

you said "1/0.01 = 68.1/x" i took it from there

OpenStudy (anonymous):

Did you multiply 0.01 * 68.1

OpenStudy (anonymous):

1/0.01 = 68.1/x 1x = (0.01)(68.1)

OpenStudy (anonymous):

ok, so it is B?

OpenStudy (anonymous):

1 kg = 1000 grams 7 kg = ? grams What did you get?

OpenStudy (anonymous):

NVM

OpenStudy (anonymous):

I don't want to give just answers...this one is easy if you show the work